Show that any prime number is either in form of 4k+1 or 4k+3, k is any positive integer


1
Expert's answer
2022-02-14T18:35:42-0500

By Euclid’s Division Lemma, each integer number "n" can be represent as "n=4k+r," where "r\in\{0,1,2,3\}."

If "r=0," then "n=4k=2(2k)," and hence "n" is a composite number.

If "r=2," then "n=4k+2=2(2k+1)," and hence "n" is a composite number.

Therefore, any prime number is either in form of "4k+1" or "4k+3," where "k" is any positive integer.